1/2< x+4
________ < or equal to 4

2
please solve

Answer by sdmmadam@yahoo.com(530)

1/2< (x+4)/2 =< 4 Multiplying by 2 (>0)
1 < (x+4) =< 8 ----(1)
(multiplication by a positive quantity does not alter the inequality)
1 < (x+4)
1-4 < x
-3 < x ----(*) [That is x >(-3)]
(x+4) =< 8
x =< 8-4
x =<4 ----(**)
Combining (*) and (**) we have -3 < x =<4
Answer: -3 < x =<4